# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ 2.2. Constructing the non-crossing $\mathrm{SLE}_{4}$ measure
A consequence of the proof is that $\mathrm{SLE}_{4}$ cannot be seen as a level-line of a single GFF, whose boundary values on the inner boundary components would not be functions of the level-line itself. Whereas we will not provide a proof, here is an intuitive argument: if a coupling $(\Phi_{D},\eta)$ of the sort were to exist, assuming that $D=\mathbb{A}_{p}$ for simplicity, we could condition on $\mathcal{A}_{+\lambda}$, the event that $\eta$ passes to the left of $C_{e^{-p}}$, and consider $D^{\prime}\subset\mathbb{A}_{p}$ a test domain passing to the left of $C_{e^{-p}}$ with $\mathrm{dist}(D^{\prime},C_{e^{-p}})>\varepsilon$. On $\mathbb{A}_{p-\delta}$, $\delta=\log\varepsilon$, we can write $\Phi_{\mathbb{A}_{p}}=\Phi_{\mathbb{A}_{p-\delta}}+u_{\xi}$ where $u_{\xi}$ is the harmonic extension of $\xi=\Phi_{\mathbb{A}_{p}}|_{C_{e^{-p}+\varepsilon}}$. The restriction property would then require that the level-line of $\Phi_{\mathbb{A}_{p-\delta}}$ conditioned to stay in $D^{\prime}$ is absolutely continuous w.r.t. the level-line of $\Phi_{D^{\prime}}^{g_{D^{\prime}}}$ with the right Radon-Nikodym derivative, and the proof shows that this is the case provided that the expectation of Equation (14) w.r.t. $\xi$ conditional on $\eta$ corresponds to $\exp(-m_{\mathbb{A}_{p}}(\eta,D^{\prime})/2)$. It is reasonable to imagine that this is only possible for all $D^{\prime}$ as above if $\xi\sim\mathcal{N}(g_{D,+\lambda},G_{D})$, so that taking $\varepsilon\to 0$ we recover that $\Phi_{D}|_{\mathcal{A}_{+\lambda}}=\Phi_{D}^{g_{D,+\lambda}}|_{\mathcal{A}_{+\lambda}}$.
The construction above determines the form of the partition function for non-crossing $\mathrm{SLE}_{4}$.
Corollary 12.
The partition function of $\mathrm{SLE}_{4}$ from $x$ to $y$ in $D$ can be expressed as
@@
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2})\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}}),$
@@
Let us explain and justify this expression. From the discussion in Remark 2.1 and the proof of Theorem 19, we may equivalently write $Z_{D}(x,y)$ as
$Z_{D}(x,y)$ $=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(\frac{1}{2}(\norm{g_{D^{\prime}}}_{\nabla(D^{\prime})}^{2}-\norm{g_{\mathbf{b}}}_{\nabla(D)}^{2}))H_{D^{\prime}}(x,y)^{1/4}\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
$=\sum_{\mathbf{b}\in\left\{\pm\lambda\right\}^{n}}\exp(2\lambda^{2}(M(\mathcal{B}^{\prime}_{-},\mathcal{B}^{\prime}_{+})-M(\mathcal{B}_{\mathbf{b}}^{-},\mathcal{B}_{\mathbf{b}}^{+})))H_{D^{\prime}}(x,y)^{1/4}\cdot\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})$
for any test domain $D^{\prime}\subset D$, without needing to regularize the Dirichlet energies. The advantage of this expression is that conformal covariance such as in Equation (7) can be read directly, given that the probability term and the one corresponding to the difference of conformal moduli are conformally invariant, as mentioned in Section 1.3.
Remark 2.3.
In [IK13, HBB10], $\mathrm{SLE}_{4}$ is studied in the annulus. For a GFF in $(\mathbb{A}_{p},e^{ix},e^{iy})$, where $x,y\in\mathbb{R}$, with boundary conditions $-\lambda$ on $(e^{iy}e^{ix})$, $+\lambda$ on $(e^{ix}e^{iy})$ and $\mu$ on $C_{e^{-p}}\subset\partial\mathbb{A}_{p}$, it is shown that the corresponding level line is driven according to the annulus Loewner equation (see [Zha04]) with a driving process $\exp(iW_{t})$, where the relative position $g_{t}(e^{ix})\exp(-iW_{t})=:\exp(iB_{t})$ is such that $B_{t}$ is a Brownian bridge from (the representative in $[0,2\pi)$ modulo $2\pi$ of) $y-x$ to $\pi(1-\mu/\lambda)$, of time-duration $p$, and up to its exit time $T$ of $[0,2\pi]$. The event that $T=+\infty$ corresponds to the resulting level-line touching the inner boundary component $C_{e^{-p}}$ (recall that the conformal modulus of the slit domain plays the role of the time parametrization), whereas $T\leq p$ corresponds to the level-line reaching $e^{iy}$ from the left or right of $C_{e^{-p}}$, when $B_{T}=0$ or $2\pi$ respectively. Observe that we have $T\leq p$ almost surely when |$\absolutevalue{\mu}\geq\lambda$, as mentioned in Section 2.1.2.
In our case, the probabilities $\mathbb{P}_{\mathbb{A}_{p}}^{g_{-\lambda}}(\mathcal{A}_{-\lambda})$ and $\mathbb{P}_{\mathbb{A}_{p}}^{g_{+\lambda}}(\mathcal{A}_{+\lambda})$ correspond to the probability that a Brownian bridge from $y-x$ to $2\pi$ (resp. to 00) exits $[0,2\pi]$ through 00 (resp. $2\pi$), and the conformal modulus of the component of $\mathbb{A}_{p}\setminus\eta_{T}$ that is doubly connected is equal to $T$.
2.3. Crossing $\mathrm{SLE}_{4}$
We now go over the construction of the $\mathrm{SLE}_{4}$ measure in the crossing case, i.e. corresponding to a measure on simple curves joining two points belonging to two different connected components of the boundary of a multiply connected domain. There are two notable differences with the non-crossing case:
  • The relevant level-lines are coupled with multivalued GFFs, corresponding to the sum of a (Dirichlet, zero-boundary) GFF and a deterministic multivalued harmonic function with locally constant boundary conditions with a change of ±$\pm 2\lambda$ around $x$ and $y$ on $\partial D$, and a change of branch describing the winding of the curve along the different boundary components. In general these level-lines are again conditioned on topological events, but in the case of doubly connected domains the coupling can be realized directly with a single compactified GFF–a sum of a (Dirichlet, zero-boundary) GFF and multivalued harmonic function with random monodromy.
  • The topological events $\mathcal{A}_{\mathbf{b}},\mathbf{b}=(b_{1},\cdots,b_{n})\in\left\{\pm\lambda\right\}^{n}$, that amounted to assignations of $-\lambda$ or $+\lambda$ as boundary conditions of a GFF to the connected components of the boundary of the domain, correspond in the crossing case to events $\mathcal{A}_{\mathbf{b}},\mathbf{b}=(b_{\text{in}},b_{1},\dots,b_{n})\in\Lambda^{n+1}:=\mathbb{Z}\times(2\lambda\mathbb{Z}+\lambda)^{n}$. They represent the change in branch of the multivalued harmonic function acting as the mean of the GFF when approaching a boundary component (target or not). The component $b_{\text{in}}$ describes the winding of any $\eta\in\mathcal{A}_{\mathbf{b}}$ around $\mathcal{B}_{\text{in}}$.
[…]
We then let $\mathcal{A}_{\mathbf{b}}$ denote the set of simple curves $\eta:x\to y$ such that the parts of $\partial U$ that are contained between any two adjacent lifts of $\eta$ correspond to constant boundary values of $g_{\mathbf{b}}$. We finally define $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ as the law of the generalized level line of a GFF $\Phi_{D}$ in $D$ with boundary conditions given by $g_{\mathbf{b}}$ (see also Appendix A).
The point of the definition is just to guarantee that $g_{\mathbf{b}}$ corresponds to the boundary conditions $\lambda(1+2k)$ on every part of the boundary of $U$ between the lifts of $\eta$ starting from $2k\pi$ and $2(k+1)\pi$ (for any $\eta\in\mathcal{A}_{\mathbf{b}}$: see Figure 4).
The Dirichlet energy $\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}$ is then defined as $\norm{g_{\mathbf{b}}}_{\nabla(D^{\prime}),\text{reg}}$, where $D^{\prime}\subset U$ is any fundamental domain for the action of the translation by $2\pi$ in $U$. For instance, one may take $D^{\prime}$ as a parallelogram with vertices $-\pi,\pi,w-\pi,w+\pi$, for any $w$ such that $e^{iw}=y$, and regularize the divergences at 00 and $w$ like in Section 1.3.
Theorem 13.
Let $D\subset\mathbb{D}$ be a crossing admissible domain. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(22) $\nu_{D}(\eta):=\sum_{\mathbf{b}\in\Lambda^{n+1}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}(\eta)\cdot Z_{D,\mathbf{b}},$
where
$Z_{D,\mathbf{b}}=\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2}),$
is the measure of $\mathrm{SLE}_{4}$ in $(D,x,y)$.
Remark 2.4.
Here again, the expression for the partition function comes from
$\frac{Z_{D,\mathbf{b}}}{Z_{D^{\prime}}}=\exp(-\frac{1}{2}\left(\norm{g_{\mathbf{b}}}_{\nabla{(D^{\prime})}}^{2}-\norm{g_{D^{\prime}}}_{\nabla{(D^{\prime})}}^{2}\right)),$
that will appear in the proof, well-defined without regularization. This also justifies conformal invariance as in the non-crossing case.
Secondly, while it is necessary to construct the measures on the events $\mathcal{A}_{\mathbf{b}}$, the restriction for these events that the winding of a curve $\eta$ around $\mathcal{B}_{\text{in}}$ is given by $b_{\text{in}}$ is actually superfluous: indeed, the level line of $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ already almost surely satisfies this property, i.e. the monodromy of $g_{\mathbf{b}}$ fixes the winding of the level-line. This was proven in [IK13] for the case of the annulus, and extends to the general case by absolute continuity of the measures: see Appendix A.
In particular, when $D$ is a conformal annulus (corresponding to $n=0$), we get the following corollary:
Corollary 14.
When $D$ is a conformal annulus, $\nu_{D}$ corresponds to the law of a level-line of a compactified GFF $\Phi=\Phi_{D}^{0}+g$, where $g$ is a random multivalued harmonic function in $D$, independent of $\Phi_{D}^{0}$, taking values $g=g_{\mathbf{b}}$ with probability $Z_{D,\mathbf{b}}/\sum_{\mathbf{b}^{\prime}\in\mathbb{Z}}Z_{D,\mathbf{b}^{\prime}}$.
These probabilities can be computed explicitly, as we will do in a second Corollary: but first, we deduce from the Theorem an expression for the partition function for crossing $\mathrm{SLE}_{\kappa}$.
Corollary 15.
The partition function of crossing $\mathrm{SLE}_{\kappa}$ can be expressed as
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\Lambda^{n+1}}\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\text{reg}}^{2})$
and it is finite. When $D$ is a conformal annulus, we have
$Z_{D}(x,y)=\sum_{b_{\text{in}}\in\mathbb{Z}}\exp(-\frac{1}{2}\norm{g_{b_{\text{in}}}}_{\nabla(D),\text{reg}}^{2}).$
@@ 2.3. Crossing $\mathrm{SLE}_{4}$
The summability of the series is provided by the exponential decay of $Z_{D,\mathbf{b}}$ as $\norm{\mathbf{b}}_{\infty}\xrightarrow[]{}\infty$. Indeed, one can write
$g_{\mathbf{b}}=g_{\left\{0,\lambda,\cdots,\lambda\right\}}+2\lambda b_{\text{in}}\cdot v_{b_{\text{in}}}+\sum_{i=1}^{n}b_{i}v_{i},$
[…]
We now turn back to computing the law of the winding of the curve under $\nu_{D}$.
Corollary 16.
If $\mathbb{P}$ denotes $\nu_{\mathbb{A}_{p}}^{\#}(x,y)$ and $\alpha$ is the representative of $(\arg y-\arg x)/(2\pi)$ in $[0,1)$, we have that
(23) $\mathbb{P}\left[\eta\in\mathcal{A}_{k}\right]=\mathbb{P}\left[\mathrm{Ind}_{\eta}(y)=k+\alpha\right]=\frac{Z_{D,k}}{Z_{D}}\varpropto e^{-\frac{\pi}{4}M(C_{e^{-p}},C_{1})\cdot(k+\alpha)^{2}}=e^{-\frac{\pi^{2}}{2p}\cdot(k+\alpha)^{2}},\quad k\in\mathbb{Z}.$
Proof.
As explained in Remark 2.4, the event $\mathcal{A}_{b_{\text{in}}}$ exactly corresponds to the event that $\mathrm{Ind}_{\eta}(y)=b_{\text{in}}+\alpha$ (with the winding defined consistently with $\alpha$), so that it suffices to compute $Z_{D,b_{\text{in}}}$ to identify the desired probabilities. Notice that $g_{b_{\text{in}}}=g_{0}+b_{\text{in}}f$, for $f=2\lambda\log\absolutevalue{z}/p$. Thus,
$\norm{g_{b_{\text{in}}}}_{\nabla(D),\text{reg}}^{2}$ $=\norm{g_{0}}_{\nabla(D),\text{reg}}^{2}+b_{\text{in}}^{2}\norm{f}_{\nabla(D)}^{2}+2b_{\text{in}}(g_{0},f)_{\nabla(D)}$
$=\norm{g_{0}}_{\nabla(D),\text{reg}}^{2}+b_{\text{in}}^{2}\cdot\frac{\pi^{2}}{p}+2b_{\text{in}}(g_{0},f)_{\nabla(D)}$
$\norm{g_{b_{\text{in}}}}_{\nabla(D),\text{reg}}^{2}=\frac{\pi^{2}}{p}\cdot(b_{\text{in}}+\alpha)^{2}+C.$
This is also $4\lambda^{2}$ times the Dirichlet energy of the harmonic function equal to $1$ on $\mathcal{B}_{\text{in}}$ and 00 on $\mathcal{B}_{\text{out}}$, and the connection with the conformal modulus is provided by Section 1.3. ∎
Note that the asymptotics for Equation (23) are consistent with taking the size of $\mathcal{B}_{\text{in}}$ to zero (or $p\to\infty$ in $\mathbb{A}_{p}$) as in this case the winding is known to be infinite almost surely [Sch00].
Proof of Theorem 13.
@@
Let $\tilde{\nu}_{D}$ be the measure in the statement of the Theorem, $D^{\prime\prime}\subset D$ be a (crossing) test domain, and let $\mathbf{b}^{\prime}\in\Lambda^{n+1}$ be such that $\left\{\eta\in D^{\prime\prime}\right\}\subset\mathcal{A}_{\mathbf{b}^{\prime}}$. This time again, instead of comparing $\tilde{\nu}_{D}$ directly to $\nu_{D^{\prime\prime}}(x,y)$, we use an intermediary domain $D^{\prime}=D\setminus\delta$, where $\delta$ is the arc of $\partial D^{\prime\prime}$ from the right of $x$ to $\mathcal{B}_{\text{in}}$. In particular, $\mathrm{SLE}_{4}$ from $x$ to $y$ in $D^{\prime}$ is the non-crossing $\mathrm{SLE}$ that was previously constructed ($D^{\prime}$ is conformally equivalent to a non-crossing admissible domain), and its law $\nu_{D^{\prime}}(x,y)$ is described by Theorem 9, in which case it holds that
@@
The set of $\mathbf{b}\in\Lambda^{n+1}$ such that $\left\{\eta\in D^{\prime}\right\}\cap\mathcal{A}_{\mathbf{b}}\neq\emptyset$ can be naturally identified with {±$\left\{\pm\lambda\right\}^{n}$, as ${\eta\in D^{\prime}}$ fixes the winding around $\mathcal{B}_{\text{in}}$, and leaves the freedom of going to the left or the right of each $\mathcal{B}_{i}$ in $D^{\prime}$. Let us under this identification consider $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$, and denote $g_{D^{\prime}}=g_{D^{\prime},\mathbf{b}},g_{D}=g_{D,\left\{b_{\text{in}},\mathbf{b}\right\}}$ (where $b_{\text{in}}$ is determined by $D^{\prime}$) for simpler notation. We have to prove that $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}$ is absolutely continuous with respect to $\mathbb{P}_{D}^{g_{D}}$ with Radon-Nikodym derivative given by
@@
This is achieved once again by computing the average of the Radon-Nikodym derivative for fixed $\xi$ over $\xi\sim\mathcal{N}(0,G_{D\setminus\eta}|_{\delta})$. Notice that conditionally on $\eta$, $\Phi_{D}^{g_{D}}|_{D^{\prime}}$ can be written $\Phi_{D\setminus\eta}^{0}+u_{\xi}+h_{\eta}$ where $u_{\xi}$ is the harmonic extension of $\xi$ to $D^{\prime}$, $h_{\eta}$ is multivalued with every branch in $D^{\prime}$ of the form $(2n-1)\lambda+2\lambda\mathbf{1}_{\text{right of }\eta}(z)$, so that the law $\mathbb{P}_{D}^{g_{D}}$ knowing $\xi$ and restricted to the event $\mathcal{A}_{\mathbf{b}}\cap\left\{\eta\subset D^{\prime}\right\}$ is the same as $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}$ from the non-crossing case thanks to the construction of $g_{D}$ and the definition of $\mathcal{A}_{\mathbf{b}}$. Similarly to the non-crossing case, we start with the computation of the Radon-Nikodym derivative for $\xi$ fixed. Although $\delta$ is one single crosscut, it is part of the boundary of both $D_{l}^{\prime}$ and $D_{r}^{\prime}$, the left and right connected components of $D^{\prime}$ relative to $\eta$: as such, we denote formally by $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$) $\delta$ seen as the boundary of $D_{l}^{\prime}$ (resp. $D_{r}^{\prime}$). We will be brought to consider Brownian excursions from $\delta$ to $\delta$ in $D$, which we understand as excursions between the two preimages of $\delta$ by $x\mapsto e^{ix}$, that we identify to $\mathcal{B}_{l}^{\prime}$ and $\mathcal{B}_{r}^{\prime}$. Denote also by $\nu_{l}$ (resp. $\nu_{r}$) the normal vector to $D_{l}^{\prime}$ (resp $D_{r}^{\prime}$) at $x\in\delta$.
Proposition 17.
@@
(24) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}}(\eta)\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\in D^{\prime}}=\mathbf{1}_{\eta\subset\mathcal{A}_{\mathbf{b}},\eta\in D^{\prime}}\exp(\frac{1}{2}(\xi,M\xi)_{\delta}-((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}},\xi)_{\delta}),$

### v2

@@ 2.2.3. Further remarks
[…]
Second, let us explain why the proof implies that the conformal restriction $\mathrm{SLE}_{4}$ cannot be seen as a level line in the usual sense where it is considered to be a local set.
Remark 2.11.
The results above show that a.s. each instance of the conformal restriction $\mathrm{SLE}_{4}$ can be seen as a sort of contour line of a single GFF: however, notice that in the construction the boundary conditions themselves depend on this contour line. As a matter of fact, a consequence of the proof is that there is no way to realise SLE as a generalised level-line of a single GFF even with random boundary conditions, as long as we do not permit the boundary conditions to depend on the lines —in particular one cannot see the conformal restriction $\mathrm{SLE}_{4}$ as a local set with respect to which the GFF would admit the usual markovian decomposition. Whereas we do not provide a proof of this claim here, let us outline an intuitive argument showing that coupling a GFF with conformal restriction $\mathrm{SLE}_{4}$ as a level line entails that the boundary conditions of the GFF are a function of the topological class of $\mathrm{SLE}_{4}$.
@@
Assume that $D=\mathbb{A}_{p}$ for simplicity and concreteness and let $D^{\prime}\subseteq D$ be a test domain with $\mathrm{dist}(D^{\prime},C_{e^{-p}})>\varepsilon$ and $C_{e^{-p}}$ belonging to the connected component of $\mathbb{A}_{p}\setminus D^{\prime}$ containing $1$. If a coupling $(\tilde{\Phi}_{\mathbb{A}_{p}},\eta)$ of the sort were to exist, we could condition on $\mathcal{A}_{+\lambda}$, the event that $\eta$ passes to the left of $C_{e^{-p}}$. On $\mathbb{A}_{p-\delta}$, $\delta$ s.t. $e^{-p+\delta}=e^{-p}+\varepsilon$, we can write $\tilde{\Phi}_{\mathbb{A}_{p}}=\Phi_{\mathbb{A}_{p-\delta}}^{0}+u_{\xi}$ where $u_{\xi}$ is the harmonic extension of $\xi=\tilde{\Phi}_{\mathbb{A}_{p}}|_{\partial\mathbb{A}_{p-\delta}}$. The restriction property would then require that the level-line of $\tilde{\Phi}_{\mathbb{A}_{p}}$ conditioned to stay in $D^{\prime}$ is absolutely continuous w.r.t. the level-line of $\Phi_{D^{\prime}}^{g_{D^{\prime}}}:=\Phi_{D^{\prime}}^{0}+g_{D^{\prime}}$ with the right Radon-Nikodym derivative, and the proof shows that this is the case provided that the expectation of Equation (18) w.r.t. $\xi$ conditional on $\eta$ corresponds to $\exp(-m_{\mathbb{A}_{p}}(\eta,D^{\prime})/2)$. It is reasonable to imagine that this property can hold for all $D^{\prime}$ as above only if $\xi\sim\mathcal{N}(g_{D,+\lambda},G_{D})$ on $C_{e^{-p}+\varepsilon}$, so that taking $\varepsilon\to 0$ we recover that on the event $\mathcal{A}_{+\lambda}$, $\tilde{\Phi}_{\mathbb{A}_{p}}|_{C_{e^{-p}}}=\Phi_{D}^{g_{D,+\lambda}}|_{C_{e^{-p}}}=+\lambda$, i.e. that the boundary condition of $\tilde{\Phi}_{D}$ on $C_{e^{-p}}$ has to be $\lambda$.
2.3. Constructing the crossing $\mathrm{SLE}_{4}$
We now go over the construction of the $\mathrm{SLE}_{4}$ measure in the crossing case, i.e. corresponding to a measure on simple curves joining two points $x,y$ belonging to two different connected components of the boundary of a multiply connected domain $D$. There are two notable differences with the non-crossing case:
  • The relevant level-lines are coupled with multivalued GFFs, corresponding to the sum of a (Dirichlet, zero-boundary) GFF and a deterministic multivalued harmonic function with locally constant boundary conditions with a change of $\pm 2\lambda$ around $x$ and $y$ on $\partial D$, and a change of branch describing the winding of the curve along the different boundary components. In general these level-lines are again conditioned on topological events, but in the case of doubly connected domains the coupling can be realised directly with a single compactified GFF–a sum of a (Dirichlet, zero-boundary) GFF and multivalued harmonic function with random monodromy.
  • The topological events $\mathcal{A}_{\mathbf{b}},\mathbf{b}=(b_{1},\cdots,b_{n})\in\left\{\pm\lambda\right\}^{n}$, that amounted to assignations of $\lambda$ or $\lambda$ as boundary conditions of a GFF to the connected components of the boundary of the domain, correspond in the crossing case to events $\mathcal{A}_{\mathbf{b}},\mathbf{b}=(b_{\text{in}},b_{1},\dots,b_{n})\in\Lambda^{n+1}:=\mathbb{Z}\times(2\lambda\mathbb{Z}+\lambda)^{n}$. The component $b_{\text{in}}$ describes the winding of any $\eta\in\mathcal{A}_{\mathbf{b}}$ around $\mathcal{B}_{\text{in}}$, while the other coordinates represent the change in branch of the multivalued harmonic function acting as the mean of the GFF when approaching the other boundary component.
We will follow a similar structure - we start with the set-up and statement, then give the proof and finish with some corollaries and remarks.
2.3.1. Set-up and statement
[…]
We then let $\mathcal{A}_{\mathbf{b}}$ denote the set of simple curves $\eta:x\to y$ such that the parts of $\partial U$ that are contained between any two adjacent lifts of $\eta$ correspond to constant boundary values of $g_{\mathbf{b}}$. We finally define $\mathbb{P}_{D}^{g_{\mathbf{b}}}$ as the law of the generalised level line of a GFF $\Phi_{D}$ in $D$ with boundary conditions given by $g_{\mathbf{b}}$ (see also Appendix A).
The point of the definition is just to guarantee that $\tilde{g}_{\mathbf{b}}$ corresponds to the boundary conditions $\lambda(1+2k)$ on every part of the boundary of $U$ between the lifts of $\eta$ starting from $2k\pi$ and $2(k+1)\pi$ (for any $\eta\in\mathcal{A}_{\mathbf{b}}$: see Figure 4).
The Dirichlet energy $\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}$ is then defined as $\norm{\tilde{g}_{\mathbf{b}}}_{\nabla(D^{\prime}),\mathrm{reg}}$, where $D^{\prime}\subset U$ is any fundamental domain for the action of the translation by $2\pi$ in $U$. For instance, one may take $D^{\prime}$ as the intersection of $U$ and a parallelogram with vertices $\pi,\pi,w-\pi,w+\pi$, for any $w$ such that $e^{iw}=y$, and regularise the divergences at 0 and $w$ like in Section 1.3.
Theorem 2.12.
Let $(D,x,y)$ be a crossing admissible domain. The measure $\nu_{D}$ on simple curves from $x$ to $y$ in $D$ defined as
(26) $\nu_{D}:=\sum_{\mathbf{b}\in\Lambda^{n+1}}\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}}}\mathbb{P}_{D}^{g_{\mathbf{b}}}\cdot Z_{D,\mathbf{b}},$
where
$Z_{D,\mathbf{b}}=\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})$
is the measure of $\mathrm{SLE}_{4}$ in $(D,x,y)$.
Corollary 2.13.
The partition function of $\mathrm{SLE}_{4}$ in a crossing admissible domain $(D,x,y)$ can be expressed as
@@ 2.3.1. Set-up and statement
$Z_{D}(x,y):=\sum_{\mathbf{b}\in\Lambda^{n+1}}\mathbb{P}_{D}^{\mathbf{b}}(\mathcal{A}_{\mathbf{b}})\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})$
@@
The summability of the series is provided by the exponential decay of $Z_{D,\mathbf{b}}$ as $\norm{\mathbf{b}}_{\infty}\xrightarrow$$\infty$. Indeed, one can write
$g_{\mathbf{b}}=g_{\left\{0,\lambda,\cdots,\lambda\right\}}+2\lambda b_{\text{in}}\cdot v_{b_{\text{in}}}+\sum_{j=1}^{n}(b_{j}-\lambda)v_{j},$
where $v_{j}$ for $1\leq i\leq n$ (resp. $v_{b_{\text{in}}}$) is a harmonic function whose boundary values are one on $\mathcal{B}_{j}$ (resp. on $\mathcal{B}_{\text{in}}$), zero elsewhere. When an index $b_{j},j\in\left\{\text{in},1,\dots,n\right\}$ goes to $\pm\infty$, the leading term in $\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2}$ is given by $b_{j}^{2}\norm{v_{j}}_{\nabla(D)}$ and therefore $\exp(-\frac{1}{2}\norm{g_{\mathbf{b}}}_{\nabla(D),\mathrm{reg}}^{2})$ decays exponentially fast.
2.3.2. Proof of the theorem
Proof of Theorem 2.12 .
@@ 2.3.2. Proof of the theorem
Let $\tilde{\nu}_{D}$ be the measure in the statement of the theorem, $D^{\prime\prime}\subset D$ be a test domain, and let $\mathbf{b}^{\prime}\in\Lambda^{n+1}$ be such that $\left\{\eta\in D^{\prime\prime}\right\}\subset\mathcal{A}_{\mathbf{b}^{\prime}}$. This time again, instead of comparing $\tilde{\nu}_{D}$ directly to $\nu_{D^{\prime\prime}}(x,y)$, we use an intermediary domain $D^{\prime}=D\setminus\delta$, where $\delta$ is the arc of $\partial D^{\prime\prime}\cap D$ from the right of $x$ to $\mathcal{B}_{\text{in}}$. In particular, $\mathrm{SLE}_{4}$ from $x$ to $y$ in $D^{\prime}$ is the non-crossing $\mathrm{SLE}$ that was previously constructed ($D^{\prime}$ is conformally equivalent to a non-crossing admissible domain), and its law $\nu_{D^{\prime}}(x,y)$ is described by Theorem 2.6, in which case it holds that
@@
The set of $\mathbf{b}\in\Lambda^{n+1}$ such that $\left\{\eta\in D^{\prime}\right\}\cap\mathcal{A}_{\mathbf{b}}\neq\emptyset$ can be naturally identified with $\left\{\pm\lambda\right\}^{n}$, as $\eta\in D^{\prime}$} fixes the winding around $\mathcal{B}_{\text{in}}$, and leaves the freedom of going to the left or the right of each $\mathcal{B}_{j}$ in $D^{\prime}$. Let us under this identification consider $\mathbf{b}\in\left\{\pm\lambda\right\}^{n}$, and denote $g_{D^{\prime}}=g_{D^{\prime},\mathbf{b}},g_{D}=g_{D,\left\{b_{\text{in}},\mathbf{b}\right\}}$ (where $b_{\text{in}}$ is determined by $D^{\prime}$) for simpler notation. We have to prove that $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}$ is absolutely continuous with respect to $\mathbb{P}_{D}^{g_{D}}$ with Radon-Nikodym derivative given by
@@
This is achieved once again by computing the average of the Radon-Nikodym derivative for fixed $\xi$ over $\xi\sim\mathcal{N}(0,G_{D\setminus\eta}|_{\delta})$. Notice that conditionally on $\eta$, $\Phi_{D}^{g_{D}}|_{D^{\prime}}$ can be written $\Phi_{D\setminus\eta}^{0}+u_{\xi}+h_{\eta}$ where $u_{\xi}$ is the harmonic extension of $\xi$ to $D^{\prime}$, $h_{\eta}$ is multivalued with every branch in $D^{\prime}$ of the form $(2k-1)\lambda+2\lambda\mathbf{1}_{\text{right of }\eta}(z)$, so that the law $\mathbb{P}_{D}^{g_{D}}$ knowing $\xi$ and restricted to the event $\mathcal{A}_{\mathbf{b}}\cap\left\{\eta\subset D^{\prime}\right\}$ is the same as $\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}$ from the non-crossing case thanks to the construction of $g_{D}$ and the definition of $\mathcal{A}_{\mathbf{b}}$. Similarly to the non-crossing case, we start with the computation of the Radon-Nikodym derivative for $\xi$ fixed. [^6] Although $\delta$ is one single crosscut, it is part of the boundary of both $D_{l}^{\prime}$ and $D_{r}^{\prime}$, the left and right connected components of $D^{\prime}$ relative to $\eta$: as such, we denote formally by $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$) $\delta$ seen as the boundary of $D_{l}^{\prime}$ (resp. $D_{r}^{\prime}$). We will be brought to consider Brownian excursions from $\delta$ to $\delta$ in $D$, which we understand as excursions between the two preimages of $\delta$ by $x\mapsto e^{ix}$, that we identify to $\mathcal{B}_{l}^{\prime}$ and $\mathcal{B}_{r}^{\prime}$. Denote also by $\nu_{l}$ (resp. $\nu_{r}$) the outwards normal vector to $D_{l}^{\prime}$ (resp $D_{r}^{\prime}$) at $x\in\delta$: we abuse notation and denote $\partial_{\nu_{l}}g_{D^{\prime}}$ for the normal derivative of $g_{D^{\prime}}$ on $\partial D_{l}^{\prime}$, and stress that $\partial_{\nu_{l}}g_{D^{\prime}}\neq-\partial_{\nu_{r}}g_{D^{\prime}}$ on $\delta$.
Proposition 2.14.
@@
(27) $\frac{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}}}}{\mathrm{d}\mathbb{P}_{D^{\prime}}^{g_{D^{\prime}},\xi}}(\eta)\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}=\mathbf{1}_{\eta\in\mathcal{A}_{\mathbf{b}},\eta\subset D^{\prime}}\exp(\frac{1}{2}(\xi,M\xi)_{\delta}-((\partial_{\nu_{l}}+\partial_{\nu_{r}})g_{D^{\prime}},\xi)_{\delta}),$